# Objective Evaluation of Prosody and Intelligibility in Speech Synthesis via Conditional Prediction of Discrete Tokens

### v1

@@ Abstract
Objective evaluation of synthesized speech is critical for advancing speech generation systems, yet existing metrics for intelligibility and prosody remain limited in scope and weakly correlated with human perception. Word Error Rate (WER) provides only a coarse text-based measure of intelligibility, while F0-RMSE and related pitch-based metrics offer a narrow, reference-dependent view of prosody. To address these limitations, we propose TTScore, a targeted and reference-free evaluation framework based on conditional prediction of discrete speech tokens. TTScore employs two sequence-to-sequence predictors conditioned on input text: TTScore-int, which measures intelligibility through content tokens, and TTScore-pro, which evaluates prosody through prosody tokens. For each synthesized utterance, the predictors compute the likelihood of the corresponding token sequences, yielding interpretable scores that capture alignment with intended linguistic content and prosodic structure. Experiments on the SOMOS, VoiceMOS, and TTSArena benchmarks demonstrate that TTScore-int and TTScore-pro provide reliable, aspect-specific evaluation and achieve stronger correlations with human judgments of overall quality than existing intelligibility and prosody-focused metrics.
@@ 1 INTRODUCTION
This paper is organized as follows: Section 2 reviews related work, highlighting the limitations of existing metrics and exploring sequence-to-sequence evaluation paradigms. Section 3 introduces the proposed metric, TTScore, detailing the text-to-speech token generator, different discrete speech tokens, and the calculation of the intelligibility and prosody score. Section 4 describes the experimental setup, datasets used, and implementation details. Section 5 presents experiments and discussions for intelligibility evaluation. Section 6 presents experiments for prosody evaluation. Finally, Section 7 concludes the paper and discusses future directions.
@@ 3 THE PROPOSED SCORE: TTScore
Motivated by this, we propose TTScore, fine-grained metrics for intelligibility, namely TTScore-int, and prosody, namely TTScore-pro, that operate directly in the speech domain. Both variants of TTScore are formulated as conditional generation tasks, where sequence-to-sequence (seq2seq) autoregressive models predict discrete speech token sequences from text. Each score is defined as the average log-likelihood of the synthesized tokens given the input text, computed with teacher forcing. TTScore-int evaluates intelligibility by measuring the alignment between content tokens and the input text, providing a finer-grained assessment than text-based metrics such as WER or CER. TTScore-pro evaluates prosody by measuring the alignment between prosody tokens and the intended content. Unlike F0-based metrics, it does not require a reference utterance; thus, it is robust to alignment errors and captures the sequential structure of prosody. Crucially, TTScore-pro can accommodate multiple valid prosodic renditions without relying on a single reference.
@@
3.A CONTENT SPEECH TOKENS
@@
3.B PROSODIC SPEECH TOKENS
For prosodic speech tokens, we adopt discrete prosody representations from FACodec [18], which has demonstrated strong prosody modeling and controllability in speech synthesis experiments with NaturalSpeech 3 [18]. FACodec prosody tokens primarily encode F0 information, one of the most crucial indicators of prosody, making them well-suited for prosody evaluation. FACodec employs residual vector quantization (RVQ) to factorize and disentangle speech into discrete representations corresponding to different attributes: prosody, content, and acoustic detail. To ensure effective disentanglement, explicit training objectives are imposed on each token type in addition to the speech reconstruction objective. For prosody tokens, these objectives include normalized F0 contour prediction and adversarial phone classification.
@@ 3.B PROSODIC SPEECH TOKENS
3.C TEXT-TO-SPEECH TOKEN GENERATOR
To estimate the distribution of discrete speech tokens in a synthesized utterance given its textual content, we train a sequence-to-sequence text-to-token generator that autoregressively predicts discrete speech tokens while incorporating content information from the input text. We adopt a transformer-based encoder–decoder architecture, specifically BART [25], which has been widely used for various sequence-to-sequence tasks. We train two text-to-token generators for TTScore-int and TTScore-pro. Both generators takes a phoneme sequence, $\mathbf{x}=[x_{1},x_{2},...,x_{L}]$, derived from the input text using a grapheme-to-phoneme converter, as the input. The difference between the two generator is the output tokens that they generate making them specialized in modeling different aspects of synthesized speech. The generator for TTScore-int predicts the content token sequence $\mathbf{c}=[c_{1},c_{2},...,c_{T}]$, for the given utterance, obtained using a pre-trained SSL model and k-means clustering as described in Section 3.A. The generator for TTScore-pro predicts phoneme-level prosody token sequence $\mathbf{f_{ph}}=[f_{1},f_{2},...,f_{L}]$ described in Section 3.B. $L$ and $T$ denote lengths of phoneme sequence and discrete speech token sequence, respectively. Both models are trained on a large-scale dataset of paired text and speech token sequences, enabling them to learn the distributions of aspect-specific tokens in natural speech.
3.D TTScore
The overall framework for the text-to-token generators and likelihood scoring is illustrated in Figure 1. Using the seq2seq text-to-token generators described in Section 3.C, with parameters $\theta_{c}$ for the content token generator and $\theta_{f}$ for the prosody token generator, we formulate the conditional probability of content tokens $\mathbf{c}$ and prosody tokens $\mathbf{f}_{\text{ph}}$ given the input phoneme sequence $\mathbf{x}$ as
@@ 3.D TTScore
$\small p(\mathbf{f_{ph}}|\theta_{f},\mathbf{x})=\prod p(f_{i}|\mathbf{f}_{<i},\mathbf{x})$ (2)
where $i\in\{1,2,...,T\}$.
This joint probability represents the expectation of the discrete speech token sequence that should occur for the corresponding content. For a given synthesized speech, if the content token sequence is likely for the given input text according to content token generator, we assume better intelligibility as the sequence contains the expected linguistic information. Similarly if the prosody token sequence is likely for the input text according to the prosody token generator, we assume a more appropriate prosody for the given linguistic information.
This probabilities can be refactored to a log-probability scores which constitues TTScore-int and TTScore-pro:
@@
$\small\text{TTScore-pro}=P(\mathbf{f_{ph}}|\mathbf{x},\theta_{f})=\frac{1}{T}\sum_{i=1}^{T}\text{log}[p(f_{i}|\mathbf{f}_{<i},\mathbf{x})]$ (4)
In order to measure intelligibility and prosody for a synthesized utterance, we first extract the corresponding discrete speech token sequences, content tokens $\mathbf{c}=[c_{1},c_{2},\ldots,c_{T}]$ and prosody tokens $\mathbf{f}_{\text{ph}}=[f_{1},f_{2},\ldots,f_{T}]$, using pre-trained speech tokenizers. The text-to-speech token generators, together with the conditional log-probability formulation, then evaluate these sequences by computing their probabilities as the average log-likelihood of each token, conditioned on the preceding tokens and the input phoneme sequence $\mathbf{x}$.
@@ 4.B MODEL ARCHITECTURE & IMPLEMENTATION DETAILS
For speech content tokens, we use the pre-trained HuBERT-base model with 12 transformer encoder layers. We experiment with discrete tokens extracted from different HuBERT layers (3, 9, 12) and with different cluster sizes (k=50, 500) for k-means. For prosody tokens, we use the pre-trained FACodec tokenizer [^2] https://github.com/lifeiteng/naturalspeech3_facodec, which operates at a 12.5 ms frame rate with a codebook size of 1024. Phoneme boundaries are obtained using the Montreal Forced Aligner [33] [^3] https://github.com/MontrealCorpusTools/Montreal-Forced-Aligner and official transcripts. Phoneme-level prosody tokens are then computed as described in Section 3.B.
@@ 4.C EVALUATION SETUP FOR INTELLIGIBILITY
To assess how well the proposed metric reflects speech intelligibility, we measure correlations between the proposed metric and standard WER/CER for synthesized speech samples. Since WER and CER are reliable and widely accepted objective metrics for intelligibility, we assume that a high correlation with them is a fundamental indicator of measuring objective intelligibility and a requirement for an objective intelligibility metric. We use Pearson linear correlation coefficient (LCC)[34] to assess the correlations on both utterance level and system level where the correlation calculated between utterance scores and averaged sytem scores, respectively. Most importantly, we evaluate how well the proposed metric reflects the perceived quality of synthesized speech by measuring correlations with human MOS ratings. In this evaluation, we compare the proposed metric with WER and CER to investigate whether measuring intelligibility in the speech domain better reflects human perception of speech. We report LCC and Spearman rank correlation coefficient (SRCC)[35] on both utterance level and system level.
@@ 4.D.3 CORRELATION WITH MOS AND TTSArena ELO
Table 1: Correlations of objective metrics with WER and CER in SOMOS and VoiceMOS datasets.
@@
SpeechLMScore HuBERT L3 50 0.08 0.22 0.11 0.21 0.32 0.42 0.31 0.40
uLM (ours) HuBERT L3 50 0.08 0.21 0.07 0.16 0.30 0.39 0.30 0.38
L3 500 0.16 0.39 0.19 0.39 0.33 0.45 0.32 0.44
L9 50 0.24 0.43 0.21 0.37 0.33 0.44 0.33 0.43
L9 500 0.42 0.69 0.43 0.68 0.44 0.55 0.45 0.57
L12 50 0.20 0.37 0.19 0.33 0.32 0.41 0.32 0.40
L12 500 0.32 0.54 0.34 0.55 0.41 0.51 0.37 0.51
TTScore-int (proposed) HuBERT L3 50 0.38 0.67 0.38 0.64 0.77 0.94 0.76 0.94
L3 500 0.37 0.70 0.43 0.71 0.75 0.93 0.74 0.96
L9 50 0.48 0.77 0.50 0.73 0.78 0.94 0.78 0.95
L9 500 0.53 0.78 0.53 0.77 0.74 0.95 0.74 0.96
L12 50 0.47 0.75 0.50 0.73 0.77 0.94 0.77 0.95
L12 500 0.48 0.76 0.53 0.77 0.77 0.96 0.77 0.97

### v2

@@ Abstract
Objective evaluation of synthesized speech is critical for advancing speech generation systems, yet existing metrics for intelligibility and prosody remain limited in scope and weakly correlated with human perception. Word Error Rate (WER) provides only a coarse text-based measure of intelligibility, while F0-RMSE and related pitch-based metrics offer a narrow, reference-dependent view of prosody. To address these limitations, we propose TTScore, a targeted and reference-free evaluation framework based on conditional prediction of discrete speech tokens. TTScore employs two sequence-to-sequence predictors conditioned on input text: TTScore-int, which measures intelligibility through content tokens, and TTScore-pro, which evaluates prosody from the perspective of pitch, through prosody tokens. For each synthesized utterance, the predictors compute the likelihood of the corresponding token sequences, yielding interpretable scores that capture alignment with intended linguistic content and prosodic structure. Experiments on the SOMOS, VoiceMOS, and TTSArena benchmarks demonstrate that TTScore-int and TTScore-pro provide reliable, aspect-specific evaluation and achieve stronger correlations with human judgments of overall quality than existing intelligibility and prosody-focused metrics.
@@ 1 INTRODUCTION
This paper is organized as follows: Section 2 reviews related work, highlighting the limitations of existing metrics and exploring sequence-to-sequence evaluation paradigms. Section 3 introduces the proposed metric, TTScore, detailing the text-to-speech token generator, different discrete speech tokens, and the calculation of the intelligibility and pitch-related prosody score. Section 4 describes the experimental setup, datasets used, and implementation details. Section 5 presents experiments and discussions for intelligibility evaluation. Section 6 presents experiments for prosody evaluation. Section 7 discusses the limitations and finally, Section 8 concludes the paper and discusses future directions.
@@ 3 THE PROPOSED SCORE: TTScore
Motivated by this, we propose TTScore, fine-grained metrics for intelligibility, namely TTScore-int, and prosody, namely TTScore-pro, that operate directly in the speech domain. Both variants of TTScore are formulated as conditional generation tasks, where sequence-to-sequence (seq2seq) autoregressive models predict discrete speech token sequences from text. Each score is defined as the average log-likelihood of the synthesized tokens given the input text, computed with teacher forcing. TTScore-int evaluates intelligibility by measuring the alignment between content tokens and the input text, providing a finer-grained assessment than text-based metrics such as WER or CER. TTScore-pro evaluates prosody by focusing on pitch patterns and measuring the alignment between prosody tokens and the intended content. Unlike F0-based metrics, it does not require a reference utterance; thus, it is robust to alignment errors and captures the sequential structure of pitch. Crucially, TTScore-pro can accommodate multiple valid prosodic renditions without relying on a single reference.
@@
3.A CHOICE OF SPEECH TOKENS
3.A.1 CONTENT SPEECH TOKENS
@@
3.A.2 PROSODIC SPEECH TOKENS
For prosodic speech tokens, we adopt discrete prosody representations from FACodec [16], which has demonstrated strong prosody modeling and controllability in speech synthesis experiments with NaturalSpeech 3 [16]. FACodec prosody tokens primarily encode F0 information, one of the most crucial indicators of prosody, making them well-suited for prosody evaluation. FACodec provides a learned, discretized, and normalized representation of pitch-related prosody, which reduces sensitivity to speaker-dependent pitch ranges, absolute F0 values, and recording conditions. FACodec employs residual vector quantization (RVQ) to factorize and disentangle speech into discrete representations corresponding to different attributes: prosody, content, and acoustic detail. To ensure effective disentanglement, explicit training objectives are imposed on each token type in addition to the speech reconstruction objective. For prosody tokens, these objectives include normalized F0 contour prediction and adversarial phone classification.
@@ 3.A.2 PROSODIC SPEECH TOKENS
3.B TTSCORE DEFINITION & METHODOLOGY
3.B.1 TEXT-TO-SPEECH TOKEN GENERATOR
To estimate the distribution of discrete speech tokens in a synthesized utterance given its textual content, we train a sequence-to-sequence text-to-token generator that autoregressively predicts discrete speech tokens while incorporating content information from the input text. Predicting specialized and distilled speech tokens is expected to be a simpler problem than speech generation itself, as the generator only needs to model limited aspects of speech in a heavily compressed discrete token space, rather than the full set of details required to reconstruct a speech waveform. This task can be even simpler for intelligibility, which can be evaluated using semantic speech tokens (e.g., HuBERT units) that already exhibit strong correspondence with input phoneme classes [28]. For this reason, text-to-specialized token generators have strong potential to effectively model token distributions due to a simpler objective, and they also tend to be more interpretable due to their focus on specific speech attributes making them suitable candidates as evaluators.
As generator, we adopt a transformer-based encoder–decoder architecture, specifically BART [19], which has been widely used for various sequence-to-sequence tasks. We train two text-to-token generators for TTScore-int and TTScore-pro. Both generators takes a phoneme sequence, $\mathbf{x}=[x_{1},x_{2},...,x_{L}]$, derived from the input text using a grapheme-to-phoneme converter, as the input. The difference between the two generator is the output tokens that they generate making them specialized in modeling different aspects of synthesized speech. The generator for TTScore-int predicts the content token sequence $\mathbf{c}=[c_{1},c_{2},...,c_{T}]$, for the given utterance, obtained using a pre-trained SSL model and k-means clustering as described in Section 3.A.1. The generator for TTScore-pro predicts phoneme-level prosody token sequence $\mathbf{f_{ph}}=[f_{1},f_{2},...,f_{L}]$ described in Section 3.A.2. $L$ and $T$ denote lengths of phoneme sequence and discrete speech token sequence, respectively. Both models are trained on a large-scale dataset of paired text and speech token sequences, enabling them to learn the distributions of aspect-specific tokens in natural speech.
3.B.2 TTScore
The overall framework for the text-to-token generators and likelihood scoring is illustrated in Figure 3. Using the seq2seq text-to-token generators described in Section 3.B.1, with parameters $\theta_{c}$ for the content token generator and $\theta_{f}$ for the prosody token generator, we formulate the conditional probability of content tokens $\mathbf{c}$ and prosody tokens $\mathbf{f}_{\text{ph}}$ given the input phoneme sequence $\mathbf{x}$ as
@@ 3.B.2 TTScore
$\small p(\mathbf{f_{ph}}|\theta_{f},\mathbf{x})=\prod p(f_{j}|\mathbf{f}_{<j},\mathbf{x})$ (2)
where $i\in\{1,2,...,T\}$ and $j\in\{1,2,...,L\}$.
This joint probability represents the expectation of the discrete speech token sequence that should occur for the corresponding content. For a given synthesized speech, if the content token sequence is likely for the given input text according to content token generator, we assume better intelligibility as the sequence contains the expected linguistic information. Similarly if the prosody token sequence is likely for the input text according to the prosody token generator, we assume a more appropriate pitch-related prosody for the given linguistic information.
These probabilities can be refactored to a log-probability scores which constitues TTScore-int and TTScore-pro:
@@
$\small\text{TTScore-pro}=P(\mathbf{f_{ph}}|\mathbf{x},\theta_{f})=\frac{1}{L}\sum_{i=1}^{L}\text{log}[p(f_{i}|\mathbf{f}_{<i},\mathbf{x})]$ (4)
In order to measure intelligibility and prosody for a synthesized utterance, we first extract the corresponding discrete speech token sequences, content tokens $\mathbf{c}=[c_{1},c_{2},\ldots,c_{T}]$ and prosody tokens $\mathbf{f}_{\text{ph}}=[f_{1},f_{2},\ldots,f_{L}]$, using pre-trained speech tokenizers. The text-to-speech token generators, together with the conditional log-probability formulation, then evaluate these sequences by computing their probabilities as the average log-likelihood of each token, conditioned on the preceding tokens and the input phoneme sequence $\mathbf{x}$.
@@ 4.B MODEL ARCHITECTURE & IMPLEMENTATION DETAILS
For speech content tokens, we use the pre-trained HuBERT-base model with 12 transformer encoder layers. We experiment with discrete tokens extracted from different HuBERT layers (3, 9, 12) and with different cluster sizes (k=50, 500) for k-means. For prosody tokens, we use the pre-trained FACodec tokenizer [^2] https://github.com/lifeiteng/naturalspeech3_facodec, which operates at a 12.5 ms frame rate with a codebook size of 1024. Phoneme boundaries are obtained using the Montreal Forced Aligner [24] [^3] https://github.com/MontrealCorpusTools/Montreal-Forced-Aligner and official transcripts. Phoneme-level prosody tokens are then computed as described in Section 3.A.2.
@@ 4.C EVALUATION SETUP FOR INTELLIGIBILITY
To assess how well the proposed metric reflects speech intelligibility, we measure correlations between the proposed metric and standard WER/CER for synthesized speech samples. Although WER and CER are imperfect, they are reliable and widely accepted objective metrics for intelligibility, we assume that a high correlation with them is a fundamental indicator of measuring objective intelligibility and a requirement for an objective intelligibility metric. We use Pearson linear correlation coefficient (LCC)[29] to assess the correlations on both utterance level and system level where the correlation calculated between utterance scores and averaged sytem scores, respectively. Most importantly, we evaluate how well the proposed metric reflects the perceived quality of synthesized speech by measuring correlations with human MOS ratings. In this evaluation, we compare the proposed metric with WER and CER to investigate whether measuring intelligibility in the speech domain better reflects human perception of speech. We report LCC and Spearman rank correlation coefficient (SRCC)[33] on both utterance level and system level.
@@ 4.D.3 CORRELATION WITH MOS AND TTSArena ELO
Table 1: Correlations of objective metrics with WER and CER in SOMOS and VoiceMOS datasets. (with bootstrapped 95% confidence intervals)
@@
SpeechLMScore HuBERT L3 50 0.08 (0.05, 0.12) 0.22 (0.09, 0.35) 0.11 (0.07, 0.15) 0.21 (0.07, 0.34) 0.32 (0.30, 0.34) 0.42 (0.28, 0.54) 0.31 (0.29, 0.33) 0.40 (0.26, 0.52)
uLM (ours) HuBERT L3 50 0.08 (0.05, 0.12) 0.21 (0.07, 0.34) 0.07 (0.03, 0.10) 0.16 (0.02, 0.29) 0.30 (0.28, 0.32) 0.39 (0.27, 0.53) 0.30 (0.28, 0.32) 0.38 (0.25, 0.52)
L3 500 0.16 (0.13, 0.20) 0.39 (0.28, 0.51) 0.19 (0.16, 0.23) 0.39 (0.28, 0.52) 0.33 (0.31, 0.35) 0.45 (0.33, 0.57) 0.32 (0.30, 0.34) 0.44 (0.32, 0.58)
L9 50 0.24 (0.21, 0.27) 0.43 (0.32, 0.55) 0.21 (0.18, 0.25) 0.37 (0.26, 0.49) 0.33 (0.31, 0.35) 0.44 (0.32, 0.57) 0.33 (0.31, 0.35) 0.43 (0.31, 0.57)
L9 500 0.42 (0.39, 0.45) 0.69 (0.63, 0.76) 0.43 (0.40, 0.46) 0.68 (0.59, 0.76) 0.44 (0.42, 0.46) 0.55 (0.45, 0.67) 0.45 (0.43, 0.47) 0.57 (0.48, 0.68)
L12 50 0.20 (0.17, 0.24) 0.37 (0.26, 0.49) 0.19 (0.15, 0.23) 0.33 (0.21, 0.45) 0.32 (0.30, 0.34) 0.41 (0.29, 0.55) 0.32 (0.30, 0.34) 0.40 (0.28, 0.54)
L12 500 0.32 (0.29, 0.35) 0.54 (0.45, 0.65) 0.34 (0.31, 0.37) 0.55 (0.46, 0.65) 0.41 (0.39, 0.43) 0.51 (0.40, 0.63) 0.37 (0.35, 0.39) 0.51 (0.40, 0.63)
TTScore-int(proposed) HuBERT L3 50 0.38 (0.35, 0.41) 0.67 (0.50, 0.76) 0.38 (0.35, 0.41) 0.64 (0.57, 0.73) 0.77 (0.76, 0.78) 0.94 (0.92, 0.96) 0.76 (0.75, 0.77) 0.94 (0.92, 0.96)
L3 500 0.37 (0.34, 0.40) 0.70 (0.64, 0.78) 0.43 (0.40, 0.46) 0.71 (0.65, 0.79) 0.75 (0.74, 0.76) 0.93 (0.91, 0.95) 0.74 (0.73, 0.75) 0.96 (0.94, 0.97)
L9 50 0.48 (0.45, 0.51) 0.77 (0.72, 0.83) 0.50 (0.47, 0.53) 0.73 (0.67, 0.80) 0.78 (0.77, 0.79) 0.94 (0.92, 0.96) 0.78 (0.77, 0.79) 0.95 (0.94, 0.97)
L9 500 0.53 (0.51, 0.56) 0.78 (0.73, 0.84) 0.53 (0.51, 0.56) 0.77 (0.72, 0.83) 0.74 (0.73, 0.76) 0.95 (0.94, 0.97) 0.74 (0.73, 0.75) 0.96 (0.95, 0.98)
L12 50 0.47 (0.44, 0.50) 0.75 (0.70, 0.82) 0.50 (0.47, 0.53) 0.73 (0.67, 0.80) 0.77 (0.76, 0.78) 0.94 (0.92, 0.96) 0.77 (0.76, 0.78) 0.95 (0.94, 0.97)
L12 500 0.48 (0.46, 0.51) 0.76 (0.71, 0.83) 0.53 (0.50, 0.56) 0.77 (0.72, 0.83) 0.77 (0.76, 0.78) 0.96 (0.95, 0.98) 0.77 (0.76, 0.78) 0.97 (0.96, 0.98)
Table 2: Correlations of objective metrics with MOS in SOMOS and VoiceMOS datasets. (with bootstrapped 95% confidence intervals)
Metric Speech token SOMOS VoiceMOS22 (all) VoiceMOS22 (test)
Utt-level Sys-level Utt-level Sys-level Utt-level Sys-level
LCC SRCC LCC SRCC LCC SRCC LCC SRCC LCC LCC
UTMOS [34] - - 0.36 (0.30, 0.41) 0.33 (0.27, 0.38) 0.53 (0.42, 0.62) 0.57 (0.46, 0.66) - - - - 0.88 (0.87, 0.89) 0.94 (0.92, 0.96)
WER Word 0.32 (0.29, 0.35) 0.30 (0.27, 0.33) 0.62 (0.54, 0.72) 0.63 (0.55, 0.72) 0.30 (0.28, 0.32) 0.31 (0.29, 0.33) 0.42 (0.30, 0.56) 0.41 (0.27, 0.56) 0.25 (0.19, 0.31) 0.39 (0.27, 0.53)
CER Character 0.34 (0.31, 0.37) 0.33 (0.30, 0.36) 0.60 (0.52, 0.70) 0.68 (0.60, 0.76) 0.28 (0.26, 0.30) 0.31 (0.29, 0.33) 0.40 (0.28, 0.54) 0.48 (0.35, 0.62) 0.23 (0.17, 0.29) 0.38 (0.25, 0.52)
SpeechLMScore [22] HuBERT (k=50) L3 0.04 (0.00, 0.08) 0.04 (0.00, 0.08) 0.12 (-0.02, 0.26) 0.08 (-0.05, 0.22) 0.48 (0.46, 0.50) 0.50 (0.48, 0.52) 0.56 (0.46, 0.68) 0.65 (0.56, 0.76) 0.47 (0.42, 0.52) 0.58 (0.48, 0.69)
uLM(ours) HuBERT (k=50) L3 0.06 (0.03, 0.09) 0.06 (0.02, 0.08) 0.17 (0.05, 0.30) 0.14 (0.01, 0.28) 0.46 (0.44, 0.48) 0.46 (0.44, 0.48) 0.53 (0.41, 0.63) 0.63 (0.52, 0.73) 0.45 (0.40, 0.50) 0.55 (0.44, 0.65)
uLM (ours) HuBERT (k=500) L3 0.17 (0.13, 0.21) 0.16 (0.12,0.20) 0.37 (0.24, 0.48) 0.36 (0.24, 0.47) 0.57 (0.55, 0.59) 0.57 (0.55, 0.59) 0.69 (0.60, 0.76) 0.72 (0.62, 0.79) 0.59 (0.55, 0.63) 0.66 (0.56, 0.74)
L9 0.34 (0.31,0.37) 0.33 (0.30, 0.36) 0.62 (0.53, 0.70) 0.62 (0.53, 0.70) 0.47 (0.45, 0.49) 0.48 (0.46, 0.50) 0.61 (0.50, 0.70) 0.63 (0.53, 0.72) 0.47 (0.42, 0.52) 0.56 (0.45, 0.66)
L12 0.24 (0.21, 0.27) 0.22 (0.18, 0.25) 0.50 (0.39, 0.60) 0.50 (0.40, 0.60) 0.44 (0.42, 0.46) 0.44 (0.42, 0.46) 0.54 (0.42, 0.64) 0.56 (0.45, 0.66) 0.43 (0.38, 0.48) 0.51 (0.38, 0.62)
TTScore-int (proposed) HuBERT (k=50) L3 0.37 (0.34, 0.40) 0.35 (0.32, 0.38) 0.64 (0.55, 0.72) 0.60 (0.49, 0.70) 0.42 (0.40, 0.44) 0.47 (0.45, 0.49) 0.48 (0.35, 0.59) 0.56 (0.44, 0.69) 0.40 (0.35, 0.45) 0.48 (0.35, 0.59)
TTScore-int (proposed) HuBERT (k=500) L3 0.40 (0.37, 0.43) 0.38 (0.35, 0.41) 0.69 (0.61, 0.76) 0.68 (0.58, 0.74) 0.48 (0.46, 0.50) 0.52 (0.50, 0.54) 0.54 (0.42, 0.64) 0.64 (0.52, 0.74) 0.47 (0.42, 0.52) 0.54 (0.42, 0.64)
L9 0.43 (0.40, 0.46) 0.42 (0.39, 0.45) 0.72 (0.64, 0.78) 0.71 (0.63, 0.78) 0.42 (0.40, 0.44) 0.46 (0.44, 0.48) 0.50 (0.37, 0.61) 0.58 (0.45, 0.69) 0.41 (0.36, 0.46) 0.51 (0.38, 0.62)
L12 0.42 (0.39, 0.45) 0.40 (0.37, 0.43) 0.71 (0.64, 0.77) 0.70 (0.62, 0.77) 0.41 (0.39, 0.43) 0.44 (0.42, 0.46) 0.48 (0.35, 0.59) 0.56 (0.44, 0.67) 0.38 (0.32, 0.44) 0.47 (0.34, 0.59)